OpenStudy (anonymous):

Convert 0.0089hL to mL i just need help setting up the problem

OpenStudy (anonymous):

1 hl = 100000 ml

OpenStudy (anonymous):

is there a way i can set it up in a fraction form?

OpenStudy (anonymous):

\[\frac{ 1 }{ 0.0089 }=\frac{ 100,000 }{ x }\]

OpenStudy (anonymous):

890ml correct?

OpenStudy (anonymous):

answer is correct does your teacher want it like this or the other way is fine \[0.0089hL \times \frac{ 100,000mL }{ 1hL } \]

OpenStudy (anonymous):

i wanted the fraction preferably for myself i dont understand the way she taught it THANKS MUCH! very appreciated! :D